OpenStudy (sherribelle):

Bare with me, fractions. If f(x) = 3x and g(x) = 1/3x, which expression could be used to verify that g(x) is the inverse of f(x)? A. 3x(x/3) B. (1/3x)(3x) C. 1/3(3x) D. 1/3(1/3x) I was thinking C but I'm not too entirely sure because the way they worded it, I don't understand it.

OpenStudy (anonymous):

g(x)=1/f(x) It is simply: 1/number

OpenStudy (sherribelle):

It's not stating 1 over f(x). It is stating one-third x.

OpenStudy (anonymous):

I think its C

OpenStudy (anonymous):

let me check one more time

OpenStudy (sherribelle):

1/3 = one-third, meaning 1/3x = one-third x.

OpenStudy (sherribelle):

So it would be one-third multiplied by 3x.

OpenStudy (anonymous):

I dont know maybe someone else can help you with this question

OpenStudy (anonymous):

@dan815

OpenStudy (sherribelle):

Thanks for trying to help.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

its C